• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Subedi

    Log entry time 18:45:36 on July 9,2010

    Entry number 331144

    keyword=Day shift summary (late entry)



    Plan for this shift (After 10 AM):

    #Continue optics data taking without sieve until about 10:45 AM.
    #Ask for escorted access to put siev in in both HRSs
    #Continue taking optics data for about 1 hr
    #At noon, ask for pass change to 2 pass. In parallel to pass change
    do the following work:
    1. Remove sieve from both HRSs
    2. Remove the small added magnet upstream of the target (Bogdan)
    3. Change polarity of the both HRSs to +ve (Tech)
    4. Change polarity of right septum to +ve (Tech)

    After the pass change, take production data.


    10:00 Optics data taking undergoing without Sieve in both HRSs.

    10:10 Right Q1 tripped.
    10:25 Asked for escorted controlled access to put sieve in.
    10:45 Access done
    10:52 Beam permit, asked for 1 uA beam on optics target with raster off.
    11:01 Now the tripped right Q1 is back.

    11:28 Target changed from optics to Tantalum.
    11:29 Asked for 2 uA beam on Tantalum
    11:50 Stopped optics data taking and asked for pass change to 2 pass.

    13:30 Sieves are out.

    14:30 Controlled access work finished, asked for beam permit. Accelerator has some issues with RF seperators.

    15:00 While turninSeptum magnet would not take any current. Asked for short controlled access so as to reset Septum magnet.

    16:00 Waiting for beam at end of shift